### v1

# HoH: A Dynamic Benchmark for Evaluating the Impact of Outdated Information on Retrieval-Augmented Generation

Abstract
While Retrieval-Augmented Generation (RAG) has emerged as an effective approach for addressing the knowledge outdating problem in Large Language Models (LLMs), it faces a critical challenge: the prevalence of outdated information in knowledge bases. Current research primarily focuses on incorporating up-to-date information, yet the impact of outdated information coexisting in retrieval sources remains inadequately addressed. To bridge this gap, we introduce HoH, the first benchmark specifically designed to evaluate the impact of outdated information on RAG. Our benchmark leverages token-level diff algorithms combined with LLM pipelines to efficiently create a large-scale QA dataset that accurately captures temporal knowledge evolution in real-world facts. Through comprehensive experiments, we reveal that outdated information significantly degrades RAG performance in two critical ways: (1) it substantially reduces response accuracy by distracting models from correct information, and (2) it can mislead models into generating potentially harmful outputs, even when current information is available. Current RAG approaches struggle with both retrieval and generation aspects when handling outdated information. These findings highlight the urgent need for innovative solutions to address the temporal challenges in RAG.
HoH: A Dynamic Benchmark for Evaluating the Impact of Outdated Information on Retrieval-Augmented Generation
Jie Ouyang, Tingyue Pan, Mingyue Chengthanks:   Corresponding author. , Ruiran Yan, Yucong Luo, Jiaying Lin, Qi Liu State Key Lab of Cognitive Intelligence, University of Science and Technology of China {ouyang_jie,pty12345,yanruiran,prime666,linjya}@mail.ustc.edu.cn {mycheng,qiliuql}@ustc.edu.cn
1 Introduction
Figure 1: An illustration of how outdated information affects RAG. The example shows a query about the US president, where the retrieved passages contain outdated information, leading to potential confusion in generating an accurate response.
Large Language Models (LLMs) have revolutionized natural language processing with their remarkable knowledge retention and reasoning abilities OpenAI et al. (2024); Grattafiori et al. (2024); Qwen (2024). However, the static nature of their pre-trained knowledge poses a significant challenge in handling rapidly evolving real-world information. LLMs frequently generate outdated responses that, while appearing plausible, no longer reflect current facts Maynez et al. (2020); Liu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022) has emerged as a promising solution to this knowledge outdating problem by dynamically retrieving and integrating up-to-date information from external sources.
Despite RAG’s promise, a critical challenge remains unaddressed: the pervasive presence of outdated information in knowledge bases Xin et al. (2024). As illustrated in Figure 1, when querying about the current US president, RAG systems may retrieve both current and outdated information, potentially leading to confusion or incorrect responses. This scenario is commonplace as information evolves and outdated facts inevitably accumulate across various sources, particularly in search engine scenarios where content is cached and redistributed. Current research primarily focuses on how to effectively retrieve and leverage the latest information Karpukhin et al. (2020); Chen et al. (2024), while overlooking how the presence of outdated information could harm RAG. This oversight is significant, as even state-of-the-art industry RAG systems Mehdi (2023) frequently generate incorrect responses due to outdated information Vu et al. (2024). To address this challenge, we introduce HoH  [^1] HoH resembles a shocked emoji [Uncaptioned image], expressing our alarm at the harm caused by outdated information. (How outdated information Harms Retrieval-Augmented Generation), the first large-scale benchmark designed to evaluate RAG’s robustness against outdated information.
HoH comprises two primary components: HoH-QA and HoH-SearchEngine. HoH-QA is a continuously evolving QA dataset that maintains temporal alignment with contemporary world knowledge. Our dataset features comprehensive evidence annotations for each QA pair, enabling fine-grained evaluation of individual RAG components. More importantly, we introduce annotations for outdated answers and evidence for the first time, allowing systematic assessment of how outdated information impacts RAG performance. Complementing the QA dataset, HoH-SearchEngine simulates real-world search scenarios by maintaining both current and historical documents, better reflecting the actual challenges RAG faced in practice. Our benchmark currently includes 96,124 QA pairs and 219,463 documents, far surpassing the scale of existing datasets.
To ensure data quality and efficiency, we develop novel methods for dataset construction and maintenance. We propose a two-stage approach that combines diff algorithms Myers (1986) with LLM pipelines to extract and verify factual changes from Wikipedia snapshots. This approach significantly outperforms previous methods Jang et al. (2022); Kim et al. (2024); Ko et al. (2024) in both efficiency and accuracy. The automated pipelines include rigorous quality control mechanisms, ensuring the reliability of the generated data while maintaining the dataset’s continuous evolution.
Through extensive experiments on the HoH benchmark, we demonstrate that outdated information significantly degrades RAG performance in two critical ways: (1) outdated information substantially reduces response accuracy: even when current information is successfully retrieved, the mere presence of outdated information in the context leads to at least 20% performance drop in mainstream LLMs, with some models performing worse than random guessing (-2.77%); (2) outdated information can mislead models into generating harmful outputs: while models maintain appropriate uncertainty when no information is retrieved ("Unknown" or "Unsure"), they become highly prone to generating confident but incorrect responses when encountering outdated information. These findings highlight the urgent need for innovative solutions to address the temporal challenges in RAG.
In summary, our main contributions are:
  1. 1.
    We introduce HoH, the first large-scale dynamic QA benchmark for evaluating RAG’s robustness against outdated information, featuring comprehensive evidence annotations and a mock search engine that maintains both current and historical documents.
  2. 2.
    We develop an efficient and scalable dataset construction method that combines traditional diff algorithms with LLMs, achieving superior quality in factual change extraction while maintaining continuous dataset evolution.
  3. 3.
    We present the first systematic analysis demonstrating the harmful impact of outdated information on RAG, providing crucial insights for future research in this domain.
Table 1: Comparison of our HoH with existing dynamic benchmarks. The automation indicates the feasibility of automatic generation. The Maintenance represents whether the validity of previously generated datasets is verified in the forthcoming time step. The evidence text indicates whether the dataset includes the evidence text. The outdated info shows whether the dataset contains annotation for outdate information. The question number and article number identify the number of the dataset question and the number of the article for retrieval, respectively.
Automation Maintenance Evidence Text Outdated Info Question Number Article Number RealtimeQA Kasai et al. (2023) 2,340 16,023 FreshQA Vu et al. (2024) 600 - TemporalWiki Jang et al. (2022) - - EvolvingQA Kim et al. (2024) 23,283 - GrowOVER Ko et al. (2024) 1,257 - PAT-Questions Meem et al. (2024) 6,172 - CLARK-News Li et al. (2024) 1,409 1,149 HoH (Ours) 96,124 219,463
2 Background and Related Work
HoH belongs to a growing body of works aimed at handling Time-Sensitive Questions in Question Answering (QA) systems. Time-Sensitive Questions are queries whose answers may vary depending on when the question is asked or the temporal context specified Chen et al. (2021). Compared to traditional QA tasks Dinan et al. (2018); Kwiatkowski et al. (2019), Time-Sensitive QA demands the effective utilization of specific temporal contexts, encompassing multiple time-evolving facts, to address time-sensitive questions Yang et al. (2024a). Based on whether temporal contexts are explicitly specified, Time-Sensitive Questions can be categorized into two types: Time-Situated Questions and Present-Anchored Questions.
2.1 Time-Situated Question Answering
Early research primarily focused on Time-Situated Questions Chen et al. (2021); Zhang and Choi (2021); Liska et al. (2022); Dhingra et al. (2022), which contain explicit temporal contexts (e.g., "Who was the President of the USA in 2023?"). Subsequent works Jia et al. (2018); Saxena et al. (2021); Tan et al. (2023) introduced more complex temporal reasoning tasks. These datasets not only contain questions with temporal contexts but also require temporal reasoning capabilities (e.g., "Who was the President of the USA before Obama?"). While these benchmarks have advanced our understanding of temporal reasoning in QA systems, they do not address the challenges posed by continuously evolving information and the need for up-to-date answers in real-world applications.
2.2 Present-Anchored Question Answering
In contrast to Time-Situated Questions, Present-Anchored Questions implicitly require the most current information available (e.g., "Who is the President of the USA?"). These questions pose unique challenges as their answers may change over time, necessitating continuous updates to maintain accuracy. Recent years have witnessed several attempts to address these challenges through dynamic benchmarks. Early approaches, such as RealtimeQA and FreshQA Kasai et al. (2023); Vu et al. (2024), rely on manual curation to ensure high-quality QA pairs, but suffer from limited scalability and impractical update frequencies. To overcome these limitations, automated approaches have emerged. TemporalWiki and EvolvingQA Jang et al. (2022); Kim et al. (2024) demonstrate the feasibility of automatic dataset generation, while GrowOVER and PAT-Questions Ko et al. (2024); Meem et al. (2024) advance this further by incorporating maintenance mechanisms. Recently, CLARK-News Li et al. (2024), in addition to providing the latest answers, is the first to record the historical progression of answer changes over time. However, due to its manual construction approach, it faces challenges in maintaining updates and is limited in scale.
As shown in Table 1, our work, HoH, builds upon these foundations while addressing several key limitations in existing benchmarks. Through automation and continuous maintenance, HoH ensures both scalability and data validity. HoH explicitly includes evidence text annotations, enabling fine-grained analyses of both retrieval and generation components in RAG systems. More importantly, it uniquely provides explicit annotations for outdated information, enabling systematic study of its impact on RAG. Lastly, HoH offers an unprecedented scale with 96,124 questions and 219,463 articles, substantially surpassing existing benchmarks in terms of data volume and comprehensiveness.
Figure 2: The overall process of HoH-QA construction comprises three primary parts: factual change extraction, QA generation, and automatic update. In the example shown, changes are identified between the July and August versions of a Wikipedia article. First, sentence pairs corresponding to factual changes are extracted. These extracted sentence pairs are then matched against previously generated datasets. For new sentence pairs, a QA generation process is conducted to create QA pairs. For sentence pairs that already exist, an automatic update process is employed to ensure the information remains current and accurate.
3 The HoH Benchmark
HoH benchmark comprises two key components: HoH-QA, a dynamic QA dataset that tracks temporal changes in facts, and HoH-SearchEngine, a mock search engine that simulates real-world information retrieval scenarios. In this section, we detail our methodology for constructing and maintaining these components.
Article Selection. HoH is based on Wikipedia snapshots  [^2] Download from https://dumps.wikimedia.org/, which contain a vast amount of world knowledge. We select almost all articles from the snapshots, excluding only low-quality articles with excessively short content (less than 200 characters). The initial version was generated based on snapshots from 2024-06-01 and 2024-07-01, with regular monthly updates since then. As of now, we have updated it to the snapshot of 2024-11-01. For more detailed statistics, see Appendix C.
3.1 HoH-QA
HoH-QA is a dynamic open-domain QA dataset that tracks temporal changes in facts through two complementary mechanisms: (1) capturing time-sensitive questions whose ground-truth answers evolve over time, and (2) maintaining continuous dataset updates to reflect the latest world knowledge. Following Jang et al. (2022); Kim et al. (2024); Ko et al. (2024), we generate HoH-QA through comparing two different Wikipedia snapshots at different time points. Building upon their approach, we enhance both the efficiency and quality, and additionally incorporate annotations for outdated information.
As illustrated in Figure 2, the main process comprises three primary parts: factual change extraction, QA generation, and automatic update.
Factual Change Extraction. The initial phase of factual change extraction is to identify the modified sections across the same article of two different snapshots. We first split the old article $A_{old}$ and new article $A_{new}$ into sentences, get $A_{old}=\{s_{old_{1}},s_{old_{2}},\cdots,s_{old_{n}}\}$ and $A_{new}=\{s_{new_{1}},s_{new_{2}},\cdots,s_{new_{m}}\}$. Following Kim et al. (2024), we employ the Myers Diff  [^3] Implemented by diff-match-patch Myers (1986) algorithm to detect modified sentence pairs. Diff algorithms are commonly employed to compare differences between texts, code (e.g., in Git), and are generally applicable to any sequence. By treating an article as a sequence of sentences, we obtain modified sentence pairs.
While prior works Jang et al. (2022); Kim et al. (2024) extract changes only at the sentence level, modified sentence pairs do not necessarily correspond to factual alterations. We further filtered these pairs and applied a diff algorithm at the character level to identify differences between sentences. For character-level differences, we employ heuristic methods to filter out modifications that are apparently not factual changes.
To further ensure the quality of extracted factual changes, we introduce a semantic screening stage using language models (LMs). Inspired by the human approach of reviewing different code versions, we consolidated the sentences before and after modification and delineated the differences between sentences in a comprehensible manner. For added parts, we use underlines for marking, and for deleted parts, we use strikethrough. To ensure that the marked differences align with the model’s reading habits, we do not process differences at the character level. Instead, we combine the diff algorithm with the tokenizer, handling differences at the token level.
In building a reliable screening model, we constructed a high-quality training dataset where three annotators independently verified 2,000 sentence pairs from our heuristically filtered samples. The resulting fine-tuned model [^4] Based on Qwen2.5-0.5B Qwen (2024) achieves 96.8% accuracy and 95.1% F1 score in identifying genuine factual changes, outperforming previous methods by a significant margin. Complete details of the fine-tuning process and comparative evaluations are provided in Appendix B.2 and A.1.
QA Generation. Following established practices in automatic QA generation Trischler et al. (2017); Rajpurkar et al. (2018), we employ LLMs to create question-answer pairs from the extracted factual changes. For each identified change, we generate a question that captures the temporal aspect, along with two distinct answer versions: the current correct answer and its outdated counterpart. The corresponding sentence pairs serve as evidence, labeled as current and outdated evidence accordingly.
The quality of generated QA pairs is ensured through a multi-stage review process. After initial generation, LLMs conduct a thorough review, focusing on the quality of both questions and answers. Detailed generation prompts and quality control procedures are provided in Appendix B.3.
Automatic Update. Our dataset maintains synchronization with evolving knowledge through monthly updates based on new Wikipedia snapshots. For each update cycle, we apply our factual change extraction pipeline to identify modifications between consecutive snapshots. The update process handles two scenarios: (1) newly emerged facts, where we directly apply our QA generation pipeline, and (2) previously captured facts that undergo further changes, which require special handling.
For facts with continuous changes, we leverage LLMs to generate updated answers to existing questions using the latest information. This approach preserves the question’s temporal relevance while creating a chain of answers that reflects the evolution of facts over time. Similar to our QA generation process, we employ LLMs to review the quality of updated answers. After confirming their accuracy and distinctness from previous versions, these answers are incorporated into our dataset, with each superseded answer preserved and marked as outdated.
3.2 HoH-SearchEngine
As QA datasets evolve over time, the corresponding external knowledge sources for retrieval must maintain temporal consistency. Search engines, being one of the most common forms of external knowledge sources, offer convenience for knowledge updates. We implement HoH-SearchEngine using Elasticsearch [^5] https://www.elastic.co/elasticsearch Elasticsearch (2018), extending its default BM25-based ranking Robertson et al. (2009) with temporal awareness. Specifically, we introduce a Gaussian decay function to discount outdated information, mimicking the temporal preference of real-world search engines. The underlying corpus comprises both current and historical versions of all articles used in the QA dataset construction, simulating the prevalence of outdated information on the web.
4 Experiment Setup
Figure 3: End-to-end performance comparison of different LLMs in the RAG evaluation. The left panel shows the overall performance metrics, while the right panel presents a detailed breakdown across categories.
In this section, we present definitions for document/passage categories, introduce time-aware instruction, explain the data and models tested in our experiments, and discuss the metrics and evaluation methods.
4.1 Type of Documents/Passages
In our study, we categorize documents/passages into three distinct types, each represented by a unique symbol, based on their relevance to the questions:
Relevant ($\mathcal{R}$): These passages contain the correct answer and are useful for the question.
Outdated ($\mathcal{O}$): These are previously relevant but have become obsolete, no longer accurate due to temporal changes.
Distracting ($\mathcal{D}$): These documents are semantically similar to the question but do not contribute correct answers.
4.2 Time-Aware Instruction and Setting
Following Vu et al. (2024), we have incorporated temporal information into the traditional RAG instruction, which is essential for addressing dynamic questions. We add the “Current Date” to the system prompt and included “Last Modified Time” for all documents, enabling temporal context awareness. Prior to response generation, the model is directed to prioritize the most current information. Figure 10 shows the complete instruction example. Additionally, we’ve implemented a time-weighted search algorithm as the default mechanism for our HoH-SearchEngine.
4.3 Data and Model Tested
Evaluation Set. We carefully curate a diverse subset of 10,000 QA samples by first clustering the entire dataset and then randomly selecting 1,000 samples from each of the 10 identified clusters.
Model Tested. We test several state-of-the-art embedding models and rerankers Chen et al. (2024); Xiao et al. (2024), including BGE-Base (bge-base-en-v1.5), BGE-M3 (bge-m3) and bge-reranker-v2-m3. Additionally, we evaluate a range of cutting-edge LLMs Qwen (2024); Grattafiori et al. (2024) spanning from 7B to 70B parameters, such as Qwen-7B (Qwen-2.5-7B-Instruct), Llama-8B (Llama-3.1-8B-Instruct), and Llama-70B (Llama-3.1-70B-Instruct).
4.4 Metrics and Evaluation
We evaluate responses using a model-based scoring system following Yang et al. (2024b). Each answer is rated by LLMs as “perfect”, “missing”, or “harmful”, scoring 1, 0, and -1, respectively. “Perfect” response accurately answers the question without hallucination. “Missing” indicates uncertainty, offering no clear benefit or detriment. “Harmful” denotes incorrect or misleading information. This method penalizes false content, prioritizing user safety by favoring omissions over errors.
5 Experiments
This section conducts a comprehensive benchmarking evaluation of various large language models and retrieval mechanisms within RAG. Through a series of experiments, our results reveal the limitations and shortcomings of existing methods when handling outdated information, providing insights for potential future enhancements.
5.1 Overview of RAG Performance
Current RAG systems demonstrate poor overall performance, with even the most capable models achieving only mediocre results. As shown on the left side of Figure 3, under our default setting, all three LLMs exhibit suboptimal performance. Even the most capable model, Llama-70B, achieves only a modest score of 51.7%. The performance degradation is more pronounced in smaller models, with Llama-8B and Qwen-7B scoring merely 40.0% and 29.9%, respectively.
The presence of outdated information significantly degrades model performance, causing not only reduced accuracy but also increased harmful outputs. To analyze this phenomenon, we categorize the retrieved information based on whether it includes relevant ($\mathcal{R}$) and outdated ($\mathcal{O}$) passages: both ($\mathcal{R}$ and $\mathcal{O}$), relevant ($\mathcal{R}$ only), outdated ($\mathcal{O}$ only) and none. As illustrated in the right side of Figure 3, when only relevant information is retrieved, all models demonstrate relatively strong performance. However, the addition of outdated information substantially reduces accuracy and increases harmful outputs.
Most concerningly, outdated information ($\mathcal{O}$) poses a greater risk than having no relevant information at all. In scenarios where only $\mathcal{O}$ is retrieved, models exhibit dangerous overconfidence, frequently generating harmful misinformation. In contrast, when no $\mathcal{R}$ or $\mathcal{O}$ is retrieved, models show appropriate uncertainty, producing fewer harmful outputs despite low accuracy. This stark contrast suggests that outdated information actively misleads models rather than merely adding noise to the generation process.
5.2 Analysis of Performance Degradation
Our previous analysis reveals that outdated information poses a fundamental challenge to RAG, leading to not only degraded performance but also increased harmful outputs. To understand the root causes of this challenge, we conduct a systematic investigation from three perspectives: (1) the retrieval module’s capability in distinguishing and filtering outdated information, (2) the generation module’s robustness against outdated information, and (3) the underlying time awareness of large language models.
Table 2: Hit rate of relevant ($\mathcal{R}$) and outdated ($\mathcal{O}$) documents for search engine results.
Decay Type @5 @10 @20 @50
- $\mathcal{R}$ 0.8707 0.9131 0.9351 0.9551
- $\mathcal{O}$ 0.8837 0.9159 0.9365 0.9747
Gauss $\mathcal{R}$ 0.7023 0.7453 0.7842 0.8294
Gauss $\mathcal{O}$ 0.4950 0.5457 0.5896 0.6463
Table 3: Hit rate of relevant ($\mathcal{R}$) and outdated ($\mathcal{O}$) passages for embedding models and reranker.
Embedding Type w/o Rerank Rerank
@5 @10 @5
bge-base-en-v1.5 $\mathcal{R}$ 0.7019 0.7365 0.7336
$\mathcal{O}$ 0.5289 0.5578 0.5563
bge-m3 $\mathcal{R}$ 0.7312 0.7578 0.7754
$\mathcal{O}$ 0.5507 0.5698 0.5678
Table 4: Performance of the generation module with passages sorted by relevance score in descending order for {×$\{\mathcal{R}\times 1,\mathcal{D}\times n\}$. Per. (perfect), Mis. (missing), Har. (harmful) and Score are in percentage.
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
$n$ Per. Mis. Har. Score Per. Mis. Har. Score Per. Mis. Har. Score
1 93.24 2.64 4.12 89.12 90.34 5.02 4.64 85.70 79.82 16.67 3.51 76.31
2 92.99 2.76 4.25 88.74 89.43 5.63 4.94 84.49 77.18 18.77 4.05 73.13
3 92.73 2.71 4.56 88.17 88.40 6.36 5.24 83.16 76.21 19.74 4.05 72.16
4 92.85 2.63 4.52 88.33 87.87 6.89 5.24 82.63 76.14 19.42 4.44 71.70
6 92.57 2.84 4.59 87.98 86.91 7.12 5.97 80.94 74.86 20.52 4.62 70.24
[…]
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
Order $n$ Per. Mis. Har. Score Per. Mis. Har. Score Per. Mis. Har. Score
[…]
1 81.13 4.74 14.13 67.00 66.87 14.44 18.69 48.18 54.46 25.65 19.89 34.57
2 80.83 4.56 14.61 66.22 68.53 12.08 19.39 49.14 53.73 26.37 19.90 33.83
3 80.44 4.17 15.39 65.05 68.45 11.55 20.00 48.45 53.90 25.58 20.52 33.38
5 79.77 4.34 15.89 63.88 69.99 9.26 20.75 49.24 54.29 24.86 20.85 33.44
[…]
1 88.39 2.67 8.94 79.45 70.15 8.74 21.11 49.04 41.42 23.85 34.73 6.69
2 87.98 2.60 9.42 78.56 68.65 7.08 24.27 44.38 40.77 23.38 35.85 4.92
3 87.12 2.70 10.18 76.94 66.09 6.13 27.78 38.31 39.77 21.79 38.44 1.33
5 86.17 2.72 11.11 75.06 63.97 4.52 31.51 32.46 37.92 21.39 40.69 -2.77
[…]
1 73.50 6.99 19.51 53.99 62.01 16.46 21.53 40.48 63.88 24.39 11.73 52.15
2 74.44 5.51 20.05 54.39 64.21 13.16 22.63 41.58 61.71 24.87 13.42 48.29
3 73.97 5.25 20.78 53.19 64.79 12.66 22.55 42.24 61.63 23.95 14.42 47.21
5 73.23 4.72 22.05 51.18 64.95 9.97 25.08 39.87 60.40 23.20 16.40 44.00
Retrieval. We first investigate the ability of the retrieval module to distinguish outdated information. Ideally, we expect relevant information ($\mathcal{R}$) to rank as high as possible while outdated information ($\mathcal{O}$) should rank lower.
Traditional search engines struggle to effectively filter out outdated information while maintaining high recall of relevant content. Our empirical investigation of search engine performance compares results with and without temporal relevance considerations (Table 2). When temporal relevance is not considered, outdated information ($\mathcal{O}$) actually shows slightly higher retrieval rates than relevant information ($\mathcal{R}$). After incorporating temporal relevance, while $\mathcal{R}$’s ranking improves relative to $\mathcal{O}$, there remains a nearly 50% probability of retrieving outdated information in the top 5 results. Moreover, this temporal awareness comes at the cost of significantly reduced recall for relevant information, with $\mathcal{R}$’s hit rate dropping by approximately 17% in top 5 results.
Current embedding models and rerankers, while effective at capturing semantic relevance, show limited ability in distinguishing temporal validity. As shown in Table 3, all tested models demonstrate strong performance in retrieving relevant information, with BGE-M3 outperforming BGE-Base, and both models showing improvements after reranking. However, these models simultaneously maintain high retrieval rates for outdated information, with hit rates consistently exceeding 50%. More problematically, models that perform better at retrieving relevant information also show proportionally better performance at retrieving outdated information, suggesting a fundamental limitation in their ability to distinguish temporal relevance.
These findings indicate an inherent conflict between maximizing relevant information retrieval and minimizing outdated content retrieval. Current mainstream methodologies appear to treat temporal relevance as a traditional relevance problem, leading to a forced trade-off between these competing objectives. This limitation at the retrieval stage places a heavy burden on the subsequent generation module to correctly identify and handle outdated information.
Table 6: Analysis of the Correlation (%) between LLMs’ Timeliness Awareness and RAG Performance.
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
$\mathcal{A_{C}}$ $\mathcal{A_{O}}$ Perfect Missing Harmful Perfect Missing Harmful Perfect Missing Harmful
24.70 2.63 72.67 33.81 15.75 50.44 26.67 23.92 49.41
52.42 11.64 35.94 52.61 23.12 24.26 44.30 33.95 21.75
76.49 0.25 23.25 68.77 7.59 23.64 61.25 16.45 22.30
93.24 2.54 4.22 83.86 8.67 7.47 76.75 16.88 6.37
Generation. Given the retrieval phase’s limitations in managing outdated information, we explore retaining such data and relying on robust generative models for its identification and processing.
[…]
Less capable models demonstrate even greater vulnerability to outdated information. The smaller models – Llama-8B and Qwen-7B – experience severe performance degradation with outdated information present. Their perfect scores fall by approximately 20%, and harmful outputs increase by up to 18%. This heightened sensitivity suggests that model capacity plays a crucial role in handling outdated information, though even the most capable models struggle significantly.
The ordering of passages critically influences model performance, with different models showing varying levels of sensitivity to passage arrangement. Table 5 reveals diverse responses to different ordering strategies across models. Llama-8B shows moderate robustness with score fluctuations within 11%, while Llama-70B maintains better overall performance but experiences variations up to 25%. Most concerningly, Qwen-7B displays extreme sensitivity with score variations exceeding 50%, sometimes performing worse than random guessing (-2.77%). These findings suggest that optimizing passage ordering could be crucial for improving RAG performance, though it cannot fully mitigate the fundamental challenge posed by outdated information.
Timeliness Awareness. The limitations observed in both retrieval and generation modules point to a more fundamental question: to what extent can LLMs understand and reason about temporal information? This investigation is crucial as it may explain the root cause of RAG’s vulnerability to outdated information. To answer this question, we conduct a systematic analysis of LLMs’ timeliness awareness capabilities.
We decompose timeliness awareness into two components: Current Awareness ($\mathcal{A_{C}}$), which measures the model’s ability to recognize current information as up-to-date, and Outdated Awareness ($\mathcal{A_{O}}$), which evaluates the model’s capability to identify outdated information as not current. These two types of awareness form the foundation for analyzing temporal understanding in LLMs.
Current LLMs demonstrate inadequate timeliness awareness, struggling to consistently differentiate between current and outdated information. As shown in Figure 4, even in an ideal scenario where models have access to both relevant and outdated information, all tested models perform poorly in timeliness awareness. While Llama-70B shows relatively balanced performance in both aspects, smaller models demonstrate clear weaknesses - Llama-8B particularly struggles with $\mathcal{A_{O}}$ (58.5%), while Qwen-7B shows poor $\mathcal{A_{O}}$ (41.3%).
Timeliness awareness capabilities strongly correlate with the performance of RAG. As shown in Table 6, models exhibit a clear performance hierarchy based on their timeliness awareness. Models possessing both types of awareness consistently achieve the best performance, while those with single awareness show moderate degradation. Most concerningly, models lacking both types of awareness perform the worst, with dramatic increases in harmful outputs. This consistent pattern suggests that enhancing LLMs’ inherent timeliness awareness could be crucial for improving the performance of RAG.
An intriguing gap exists between outdated awareness and harmful output prevention. Notably, some models with only $\mathcal{A_{O}}$ produce more harmful outputs than those with only $\mathcal{A_{C}}$, despite being able to recognize outdated information. Llama-70B demonstrates this paradox strikingly, generating 11% more harmful outputs with $\mathcal{A_{O}}$ compared to $\mathcal{A_{C}}$. This suggests that merely identifying information as outdated does not guarantee the model will avoid using it harmfully. Further alignment might be needed to ensure models refrain from generating harmful outputs when they recognize the information as outdated.
Figure 4: Timeliness awareness (%) of LLMs ideally, where retrieval results always contain both $\mathcal{R}$ and $\mathcal{O}$.
6 Conclusion
We introduced HoH, the first benchmark designed to evaluate how outdated information impacts RAG systems. By novelly combining token-level diff algorithms and language models, we efficiently constructed a high-quality large-scale QA dataset. Our experiments demonstrated that outdated information significantly degrades RAG performance by reducing accuracy and potentially generating harmful outputs. Current RAG approaches struggle with both retrieval and generation aspects when handling outdated information. This work provides a new perspective on RAG vulnerabilities and offers crucial infrastructure for future improvements.
7 Limitation
We use changes in Wikipedia snapshots to generate and update benchmarks, which effectively capture the evolution of real-world knowledge but also have several limitations. Firstly, Wikipedia snapshots have a fixed update frequency, while real-world knowledge changes continuously at varying speeds. For rapidly evolving fields like the stock market, our benchmarks cannot reflect real-time changes. Secondly, our QA datasets are primarily derived from individual articles, which may limit their effectiveness for tasks that require synthesizing information from multiple sources. Lastly, despite our efforts to closely simulate a real-world search engine, our mock search engine has significant limitations. This is mainly due to the homogenization of outdated information, as our outdated corpora primarily come from the historical versions of the same article, making them very similar to each other. In contrast, real-world outdated information is diverse, often sourced from various origins, and displays significant differences.
Appendix A Effectiveness and Efficiency
A.1 Effectiveness of Change Extraction
Accurate extraction of factual changes between different versions of Wikipedia snapshots is crucial for building high-quality dynamic QA datasets. Previous works have adopted various approaches to address this challenge. EvolvingQA Kim et al. (2024) relied primarily on traditional diff algorithms to identify changes, while GrowOVER Ko et al. (2024) employed a similarity-based approach using sentence embeddings generated by SimCSE Gao et al. (2021).
EvolvingQA. This method uses character-level diff algorithms to detect modifications between text versions, which directly compares the textual differences at character level.
GrowOver. This approach adopts a dual-tower architecture where sentence pairs are independently encoded using SimCSE embeddings. The cosine similarity between these embeddings is then used to determine if the sentences contain substantial changes.
[…]
HoH. Our full model employs token-level diff information in the input structure, which aligns better with the token-based nature of language models.
As shown in Table 7, previous methods achieve relatively low performance in this task. The diff-based approach used in EvolvingQA only achieves 20.85% accuracy and 34.50% F1 score, as it lacks semantic understanding and treats all textual changes equally. The embedding similarity method performs slightly better but still only reaches about 41% on both metrics.
[…]
These results validate the effectiveness of our approach in identifying factual changes, which is fundamental to ensuring the quality of our HoH dataset.
A.2 Efficiency Analysis
Given the computational costs and complexity involved in dynamic dataset construction, we provide a theoretical analysis of the efficiency of different approaches. The construction process primarily consists of two key steps: (1) factual change extraction and (2) QA generation/update using LLMs. We analyze the time complexity for processing a single article:
Let $T_{ext}$ denotes the average time for extracting sentence pairs from an article, $N$ be the average number of sentence pairs extracted per article, and $T_{llm}$ represents the processing time for LLM per sentence pair. The total processing time $T_{total}$ for an article can be expressed as:
$T_{total}=T_{ext}+N\times T_{llm}$
As shown in Table 8, the traditional sentence-level diff algorithm (as used in EvolvingQA) initially extracts a large number of sentence pairs (over 3.7M pairs across five months). While $T_{ext}$ is relatively small with this approach, $N$ remains large as it captures all sentence-level changes without distinguishing factual from non-factual modifications. The embedding-based approach (GrowOVER) requires computing and comparing embeddings for all sentence pairs, increasing $T_{ext}$ while still struggling to effectively reduce $N$.
Our method achieves better efficiency through a multi-stage filtering process: (1) Initial heuristic filtering reduces the number of sentence pairs by approximately 74.57%, eliminating obvious non-factual changes. Moreover, (2) language model screening further filters out 85.80% of the remaining pairs, identifying subtle non-factual modifications. Therefore, (3) the final LLM processing only needs to handle about 133,973 sentence pairs, significantly reducing the computational cost of the most time-consuming component ($T_{llm}$)
This empirical evidence supports our theoretical analysis that our approach achieves better computational efficiency by effectively reducing $N$ through accurate filtering, even though we may spend slightly more time on initial change detection ($T_{ext}$). The dramatic reduction in sentence pairs (from 3.7M to 133K) before LLM processing demonstrates the effectiveness of our filtering strategy in minimizing unnecessary computational costs.
Appendix B Details of Benchmark Construction
B.1 Details of Heuristic Filtering
We employ heuristic methods to filter out the modified sentence pairs which are obviously not factual changes. The modifications mainly include the following categories:
  • Pronoun Changes: Changes in pronouns, such as “he” to “James”. These changes typically do not affect the factual content of the sentence but merely adjust the perspective or grammatical structure.
  • Spelling Corrections: Corrections of misspellings(edits with a distance less than 2, excluding numerical changes).
  • Frequent Changes: Statistically frequent modifications, such as “USA” to “United States”. These might result from common stylistic differences, such as phrase replacements or grammatical adjustments.
If the sentence pair differences only involve the above types of modifications, or solely additions or deletions, we simply remove them. With these operations, we substantially reduce the volume of data for subsequent processing. Meanwhile, through heuristic filtering, we eliminate some noisy data, further ensuring data quality when processed by the language model.
B.2 Details of Fine-Tuning Methods
Despite the fact that the Heuristic Filter can effectively filter out most sentence pairs that clearly do not exhibit factual changes, the number of remaining sentence pairs is still substantial. Directly using a large language model (LLM) for analyzing and generating questions based on these sentence pairs would result in significant resource consumption. Therefore, we train a language model-based sentence pair discriminator to further filter out sentence pairs without factual changes at the semantic level.
Specifically, we randomly select approximately 2000 sentence pairs from a dataset that had undergone heuristic filtering, and these pairs are annotated and mutually verified by three annotators. Among these sentence pairs, about 400 pairs that exhibited factual changes are labeled as 1, while the remaining over 1600 pairs without factual changes are labeled as 0. Subsequently, we use Qwen2.5-0.5B Qwen (2024) as the backbone model and replace its output layer with a linear classification head. We split the dataset into training and test sets with a ratio of 8:2, then perform full fine-tuning of Qwen-0.5B on the training set, with the objective of minimizing cross-entropy loss, and evaluate on the test set. This experiment is conducted using PyTorch Paszke et al. (2019) on a single NVIDIA A100 40G GPU, with a learning rate set to 2e-5 and a batch size of 24; other parameters can be found in the config file of our open-source repository.
B.3 Details of QA Generation
We use Llama-3.1-70B-Instruct for QA generation. As shown in the prompt in Figure 5, we instruct it to identify the contradiction between sentence pairs and generate a QA pair that reflects this contradiction. The question should be answerable based on each sentence, while ensuring the two answers contradict each other. In addition to the sentence pairs, we also provide broader context from the original text containing these sentences. However, we specify that the QA pair should relate directly to the information presented in the old/new sentence(s).
After generating questions, we perform a comprehensive quality review. The quality review contains the following parts:
  • Same Answer Check: The two generated answers must exhibit semantic discrepancies, ensuring meaningful and distinct differences between the original and updated answers.
  • Answer Accuracy: Answers must strictly align with the provided context, avoiding errors or irrelevant information, and fully capture key details.
  • Question Clarity: Questions should be precisely formulated with explicit terminology, avoiding ambiguity or unclear references.
  • Question Completeness: Questions must be self-contained, providing all necessary details without requiring external context.
  • Temporal Independence: Questions should avoid references to specific times or updates, focusing solely on the content itself.
Prompts are shown in Figure 6 and Figure 7. Only when the quality review is fully passed, we incorporate the generated QA pairs into the dataset; otherwise, we require the LLM to regenerate them. If the maximum number of regeneration attempts (3) is reached, we discard them.
B.4 Details of Automatic Update
We omit how to distinguish between newly emerged sentence pairs and pre-existing sentence pairs in the main text. Taking sentence pair $S=(s_{old},s_{new})$ of article $A$ as an example. We first identify all QA pairs generated from article $A$ in the dataset and locate all corresponding evidence $E$. If the knowledge in the dataset has changed again, then there should be some evidence $e_{i}\in E$ that matches $s_{old}$. If any $e_{i}\in E$ matches $s_{old}$, we identify $S$ as a pre-existing sentence pair. For pre-existing sentence pairs, we subsequently employ prompt shown in Figure 8 to generate a new answer. The quality review prompt is the same as QA Generation.
Appendix C Dataset Statistics
C.1 Field Description
Table 17 shows the detailed field description of HoH-QA. Each entry in our dataset contains a question, its current valid answer with supporting evidence, and a document reference. Notably, through the outdated_infos field, we also track the historical evolution of answers, maintaining a chronological record of previous valid answers along with their corresponding evidence and modification dates. Table 18 provides several concrete examples illustrating these different fields and how they capture the temporal progression of answer updates.
C.2 Statistic
To help users better understand our dataset, we conduct a statistical analysis of its contents. Table 9 summarizes the number of new questions generated in different months, as well as the number of questions that require updated answers, while Table 10 shows the distribution of questions based on how many times their answers have been updated. Furthermore, Table 11 summarizes the number of questions from various domains in the dataset, which is consistent with the iteration rates of these domains in the real world.
Table 9: Statistic of QA generation in different months, where "New" indicates the number of new questions generated in the month, while "Updated" refers to the number of questions that require updated answers during the same period.
Month New Updated
06-07 16896 -
07-08 20453 683
08-09 20047 1112
09-10 19466 1506
10-11 19262 1828
Total 96124 5129
Table 10: Distribution of questions by the number of outdated answers, where the number also indicates how many times the answer has been updated.
Number 1 2 3 4 5
Count 91,856 3,647 447 108 66
Table 11: Distribution of questions across domains.
Class Number Percentage
Finance 22,786 23.70%
Sports 30,011 31.22%
Music 5,592 5.82%
Movie 6,038 6.28%
Open 31,697 32.98%
Total 96,124 100%
Appendix D Additional Experiment Setup
D.1 Output Type
In QA tasks, two predominant output modes exist: long answer and short answer Kwiatkowski et al. (2019). Long answers consist of detailed information and longer text passages. Such answers may comprise a paragraph or multiple sentences, aiming to provide comprehensive information in response to questions. Short answers usually consist of a few words, phrases, or sentences that directly address the core content of the question, emphasizing precision and directness. While long answers are more conducive to human comprehension and align better with LLM’s natural output patterns, short answers are more commonly used in evaluations due to the challenges in assessing long-form responses.
In this study, we evaluate both output modes. For the long answer (Figure 13), we impose no restrictions on the LLM’s output, allowing it to generate responses freely. For the short answer (Figure 10), we constrain the LLM to provide minimalist responses, with an explicit instruction to indicate uncertainty through an “Unsure” response. Due to space constraints, the experimental results for the long answer are not presented in the main text.
D.2 Retrieval Metrics
We use top-$k$ ($k\in\{5,10,20,50\}$) hit rate as the metric for retrieval.
Suppose we have a list of retrieval results $\{r_{1},r_{2},\ldots,r_{n}\}$, where each $r_{i}$ is a retrieved item. The binary Hit Rate is computed using the following formula:
$HR@k=\left\{\begin{array}[]{ll}1,&\text{if any }r_{i}\in\mathcal{R}\text{ or }\mathcal{O},\\ 0,&\text{otherwise.}\end{array}\right.$
This metric is calculated separately for relevant and outdated information to compare their relative rankings in the retrieval results.
D.3 Model-based Auto Evaluation
Similar to previous work Xu et al. (2023); Yang et al. (2024b), we employ model-based automatic evaluation, using LLMs as judges. In this experiment, we utilize Llama-3.1-70B-Instruct as the judge. However, a fundamental prerequisite for employing LLMs as judges is that they must exhibit sufficient accuracy. To verify this, we conduct a manual evaluation of Llama-3.1-70B-Instruct’s assessment results. As shown in Table 12, the performance of model-based automatic evaluation approaches perfection.
Table 12: Manual assessment of model-based automatic evaluation. All numbers are in percentage.
Precision Recall F1 Score
Perfect 98.5 100.0 99.3
Harmful 100.0 98.5 99.2
Missing 100.0 100.0 100.0
D.4 Default End-to-end Setting
End-to-end Experiments. Our experiments follow the standard RAG solution. Upon receiving a user query, the system initially retrieves $k$ documents from the internet deemed most relevant to the query via a search engine Lazaridou et al. (2022). From these $k$ documents, the retriever attempts to find a sufficiently small subset that enables the generator to accurately answer the query Karpukhin et al. (2020). First, we split each document into passages. When splitting, we maintain passage length within $c_{s}$ tokens and allow $c_{o}$ tokens overlap between consecutive passages while preserving sentence and paragraph integrity. These passages serve as the basic units for retrieval. If reranking is applied, we first retrieve the top $m$ passages using embedding models and then further rerank these passages to obtain the top $n$ passages. If reranking is not applied, we directly retrieve the top $n$ passages using embedding models. The generator, typically a LLM, subsequently synthesizes these $n$ relevant passages to produce the final answer to the query. By default, the system employs bge-m3 without reranking, utilizing the following parameters: $k=20,c_{s}=256,c_{o}=32,n=5$.
Generation Experiments. Ideally, we assume the retrieval step consistently retrieves relevant passages. Keeping other parts of the LLM input constant, we only modify the passage type and quantity. Symbolic notation, such as {×$\{\mathcal{R}\times 1,\mathcal{O}\times 1,\mathcal{D}\times 2\}$, denotes 1 relevant passage ($\mathcal{R}$), 1 outdated passage ($\mathcal{O}$), and 2 distracting passages ($\mathcal{D}$). Since for most questions, $\mathcal{R}$ and $\mathcal{O}$ typically have only one instance, we primarily conduct experiments by adjusting the number of $\mathcal{D}$. Distracting passages are selected based on descending retrieval scores, aligning with standard RAG settings.
D.5 Detailed Setup
Model Setting. Due to budget constraints, we do not test any close-source LLMs. For all the open-source LLMs tested (Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct), we locally deploy them in bfloat16 using vLLM  [^6] We deploy a OpenAI Compatible Server using vLLM. on our own server. The temperature parameter for each model is set to 0.3 to reduce output uncertainty, and the maximum output length is limited to 100 tokens.
Environment Setting. All experiments are completed on a Linux server with AMD EPYC 7742 64-Core Processor CPUs @ 2.25GHz and 8 NVIDIA A100 GPUs (40G). GPUs are used for deploying close-source models. The version of Python is 3.10.15. The version of the torch package is 2.5.1. The version of the transformers package is 4.46.2. The version of the vllm package is 0.6.4.post1.
Appendix E Additional Experiments
E.1 Vanilla Results
Table 13: Performance of the generation module without retrieval. Per.(perfect), Mis.(missing), Har.(Harmful) and Score are in percentage.
Model Per. Mis. Har. Score
Llama 3.1 70B 9.38 62.03 28.59 -19.21
Llama 3.1 8B 4.09 80.84 15.07 -10.98
Qwen 2.5 7B 0.66 97.21 2.13 -1.47
In addition to the RAG experiments, we also conduct LLM-only experiments. As shown in Table 13, for these dynamic questions, when no external resources are available, all LLMs demonstrate considerable caution, declining to answer the majority of questions.
E.2 Additional Results for Generator
We conduct additional experiments on the effect of passage types on generation performance. As shown in Table 14, when only outdated passages($\mathcal{O}$) are available without relevant $\mathcal{R}$), the generator’s performance indicate that LLMs are inevitably misled by outdated information, producing a high proportion of potentially harmful responses.
Moreover, we conduct additional experiments focusing on long answers. As demonstrated in Table 15 and Table 16, the experimental findings largely align with those observed in short-answer scenarios. The impact of distracting passages($\mathcal{D}$) remains substantially lower than that of outdated passages($\mathcal{O}$), and all three models exhibit significant sensitivity to passage ordering. One notable difference is that Qwen-7B demonstrates marked improvement in long-answer performance, with its optimal performance even surpassing Llama-8B under appropriate passage sequencing conditions.
Table 14: Performance of the generation module with passages sorted by relevance score in descending order for $\{\mathcal{O}\times 1,\mathcal{D}\times n\}$. Per. (perfect), Mis. (missing), Har. (harmful) and Score are in percentage.
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
$n$ Per. Mis. Har. Score Per. Mis. Har. Score Per. Mis. Har. Score
1 11.92 5.01 83.07 -71.15 10.96 9.31 79.73 -68.77 8.21 20.58 71.21 -63.00
2 13.96 5.01 81.03 -67.07 12.41 10.27 77.32 -64.91 8.30 22.69 69.01 -60.71
3 14.82 4.98 80.20 -65.38 13.31 10.59 76.10 -62.79 8.59 23.73 67.68 -59.09
4 15.29 4.79 79.92 -64.63 13.76 11.54 74.70 -60.94 8.67 23.04 68.29 -59.62
6 15.82 4.72 79.46 -63.64 14.22 11.37 74.41 -60.19 9.04 24.02 66.94 -57.90
Table 15: Long Answer performance of the generation module with passages sorted by relevance score in descending order for {×$\{\mathcal{R}\times 1,\mathcal{D}\times n\}$. Per. (perfect), Mis. (missing), Har. (harmful) and Score are in percentage.
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
$n$ Per. Mis. Har. Score Per. Mis. Har. Score Per. Mis. Har. Score
1 92.19 3.12 4.69 87.50 91.09 4.14 4.77 86.32 90.28 2.35 7.37 82.91
2 91.79 3.01 5.20 86.59 89.95 4.48 5.57 84.38 89.49 2.52 7.99 81.50
3 91.60 3.12 5.28 86.32 89.01 4.85 6.14 82.87 89.19 2.78 8.03 81.16
4 91.54 3.26 5.20 86.34 88.99 4.63 6.38 82.61 88.49 2.89 8.62 79.87
6 91.36 3.24 5.40 85.96 88.58 4.61 6.81 81.77 88.13 3.05 8.82 79.31
[…]
Llama 3.1 70B Llama 3.1 8B Qwen 2.5 7B
Order $n$ Per. Mis. Har. Score Per. Mis. Har. Score Per. Mis. Har. Score
[…]
1 77.84 8.17 13.99 63.85 68.35 13.21 18.44 49.91 66.14 3.98 29.88 36.26
2 75.61 9.55 14.84 60.77 69.21 11.85 18.94 50.27 66.59 4.31 29.10 37.49
3 75.04 9.91 15.05 59.99 68.20 12.04 19.76 48.44 66.42 4.39 29.19 37.23
5 73.92 11.02 15.06 58.86 68.28 11.05 20.67 47.61 66.53 4.92 28.55 37.98
[…]
1 83.43 6.36 10.21 73.22 68.72 10.95 20.33 48.39 48.20 5.28 46.52 1.68
2 81.91 7.21 10.88 71.03 65.72 10.71 23.57 42.15 48.36 5.45 46.19 2.17
3 80.85 7.70 11.45 69.40 63.28 10.51 26.21 37.07 46.08 5.60 48.32 -2.24
5 79.37 7.93 12.70 66.67 61.53 9.57 28.90 32.63 44.75 6.38 48.87 -4.12
[…]
1 70.33 11.22 18.45 51.88 66.55 13.83 19.62 46.93 77.83 2.99 19.18 58.65
2 68.72 12.13 19.15 49.57 66.31 12.70 20.99 45.32 75.59 3.32 21.09 54.50
3 67.86 12.60 19.54 48.32 64.95 12.91 22.14 42.81 75.57 3.39 21.04 54.53
5 66.65 12.88 20.47 46.18 63.25 12.01 24.74 38.51 73.91 3.63 22.46 51.45
Table 17: Field Description
Field Type Element Description
question str —- The question for RAG to answer.
answer str —- The latest gold answer to the question.
evidence str —- Source or reference material that substantiates the answer.
last_modified_time date —- The last modified date of the document containing the answer.
outdated_infos list[dict] answer A collection of previously outdated valid responses, each entry with its answer, evidence, and modification date.
evidence
last_modified_time
document dict id Information about the source document, including a unique identifier and its title.
title
Table 18: Examples of QA
Example 1
Document:
         id: 100011
         title: Otago
Question: What was the median income in Otago compared to the national median income?
Answer: \$39,100, compared with \$41,500 nationally
Evidence: The median income was \$39,100, compared with \$41,500 nationally.
Last modified time: 2024-11-01
Outdated_infos:
         [Info 1]:
                Answer: \$30,000, compared with \$31,800 nationally
                Evidence: The median income was \$30,000, compared with \$31,800 nationally.
                Last modified time: 2024-10-01
Example 2
Document:
         id: 1000214
         title: Master of the Horse
Question: Who is the current Master of the Horse?
Answer: The Lord Ashton of Hyde
Evidence: The current Master of the Horse is The Lord Ashton of Hyde.
Last modified time: 2024-07-01
Outdated_infos:
         [Info 1]:
                Answer: Lord de Mauley
                Evidence: The current Master of the Horse is Lord de Mauley.
                Last modified time: 2024-06-01
Example 3
Document:
         id: 1014556
         title: Jamaica national football team
Question: What is the opponent in the match after which the caps and goals for the Jamaica national
football team are correct?
Answer: Venezuela
Evidence: Caps and goals correct as of 30 June 2024, after the match against Venezuela.
Last modified time: 2024-08-01
Outdated_infos:
         [Info 1]:
                Answer: Dominica
                Evidence: Caps and goals correct as of 9 June 2024, after the match against Dominica.
                Last modified time: 2024-07-01
         [Info 2]:
                Answer: Canada
                Evidence: Caps and goals correct as of 21 November 2023, after the match against Canada.
                Last modified time: 2024-06-01
Figure 5: Sample prompt for QA Generation.
Figure 6: Sample prompt to evaluate the quality of a generated QA pair. Above is the system prompt defining the evaluation criteria for the quality of a generated QA pair. Below is the user prompt along with the QA pair generated for evaluation purposes.
Figure 7: Sample prompt to check if the old answer and the new answer are semantically the same. The interior box provides a specific example, including a question, two answers, and the final evaluation response.
Figure 8: Sample prompt for new answer generation. The interior box offers a specific example, including an old question, sentences, source content, and the old answer.
Figure 9: Sample prompt for model-based evaluation.
Figure 10: Sample prompt for RAG with Time-Aware Instruction (Short Answer). This prompt emphasizes providing up-to-date information based on document timestamps. The upper section shows the system prompt with current dates noted for context.
Figure 11: Sample prompt for evaluate LLMs’ timeliness awareness.
Figure 12: Sample prompt for generation without retrieval (Vanilla).
Figure 13: Sample prompt for RAG with Time-Aware Instruction (Long Answer). This prompt emphasizes providing up-to-date information based on document timestamps. The upper section shows the system prompt with current dates noted for context.